f(x) = √(cosx) + √(sinx) dx
1
Expert's answer
2020-06-01T18:05:06-0400
"f(x)=\sqrt{(\cos x)}+\sqrt{(\sin x)}"

"{df\over dx}={1\over 2\sqrt{(\cos x)}}(-\sin x)+{1\over 2\sqrt{(\sin x)}}(\cos x)="

"={\cos x\over 2\sqrt{(\sin x)}}-{\sin x\over 2\sqrt{(\cos x)}}"